OpenStudy (anonymous):

Waterslide fun! (not really).. A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 32 feet and the child ride has a radius of 24 feet, what is the ratio between the volumes of each ride? Answer 4:3 16:9 64:27 12:9

OpenStudy (pfenn1):

Do you know how to calculate the volume of a cone?

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

So the ratio of the radii of the adult ride to the child's ride must be equal to the ratio of the height of the adult ride to the child's ride. The ratio of radii = 32:24=4:3. But they are asking for the ratio of the Volumes of each ride. What do you suppose that ratio would be?

OpenStudy (anonymous):

hmm.. im not sure.. would we figure out the volume of both?

OpenStudy (pfenn1):

We don't have enough info to figure out the volume of both, we don't know the height of either. But we know that the ratio of the radii has to be the ratio of the heights as well (because the cones are similar). So the height of the adult ride is 4/3 the height of the child ride or 4/3 h.

OpenStudy (anonymous):

ok.. so now what lol..

OpenStudy (pfenn1):

Let r=radius of kid ride, then radius of adult ride = (4/3)r Let h=height of kid ride, then height of adult ride = (4/3)h The volume of a cone is given by\[V=\frac13\pi r^2h\]So what is the ratio of the volume of the adult ride to the volume of the kid ride?

OpenStudy (anonymous):

A?

OpenStudy (pfenn1):

You need to evaluate the volume of the adult over the volume of the kid \[\frac{V _{a}}{V _{k}}=\frac{\frac13\pi r _{a}^2h _{a}}{\frac13\pi r _{k}^2h _{k}}=\frac{ r _{a}^2h _{a}}{ r _{k}^2h _{k}}\]but the (adult r) = (4/3)r and (adult h) = (4/3)h so\[=\frac{(\frac 43 r)^2h}{\frac 43 r^2h}\] Can you simplify this?

OpenStudy (anonymous):

no?

OpenStudy (pfenn1):

Sorry. That last expression is wrong. It should have been\[=\frac{(\frac 43 r)^2\frac43 h}{\frac 43 r^2h}\] Can you simplify this?

OpenStudy (anonymous):

i don't see a way to simplify..

OpenStudy (pfenn1):

Would it be easier if I presented it like this?\[=\left( \frac{(\frac43)^2\frac43}{\frac43} \right)\left( \frac{r^2}{r^2} \right)\left( \frac hh \right)\]

OpenStudy (anonymous):

eh.. i still wouldn't know what to do.

OpenStudy (pfenn1):

What is the simplification of\[2\over 2\]

OpenStudy (pfenn1):

or how about \[x^2 \over x\]

OpenStudy (anonymous):

1?

OpenStudy (pfenn1):

\[\frac 22=1\]\[\frac{x^2}{x}=x\]

OpenStudy (pfenn1):

So what is \[\frac{r^2}{r^2}=?\]\[\frac hh=?\]

OpenStudy (pfenn1):

And if you say 4/3=x, then what is \[\frac{x^3}{x}=?\]

OpenStudy (anonymous):

4/3 h r^2

OpenStudy (pfenn1):

Not quite.\[\left( \frac{(\frac43)^2\cancel{\frac43}}{\cancel{\frac43}} \right)\left(\cancel{ \frac{r^2}{r^2}} \right)\left( \cancel{\frac hh} \right)=\left( \frac43 \right)^2\left( 1 \right)\left( 1 \right)=\frac{16}{9}\]

OpenStudy (anonymous):

Oh. I just don't really understand it.. but thanks for your help.

OpenStudy (pfenn1):

Well, I tried to explain it but......hmmmm. Sorry.

OpenStudy (anonymous):

np :/